# Deciphering the global production network from cross-border firm transactions

### v2

@@ Methods
This paper relies on data provided by the UK HM Government’s Global Supply Chain Intelligence Programme (GSCIP). This contains firm-to-firm cross-border trade, derived from customs and shipping data, with details on the transacting firms compiled from tax filings, company websites, and other sources. Firm-to-firm transaction data is derived from administrative customs data. GSCIP amalgamates first-party customs data across a multitude of countries.
Our dataset includes transaction information for every country. While the dataset does not have first-hand customs data or another primary source of transaction information for every country, the data that we do have includes transactions between firms within the countries included and those out of sample providing partial coverage.
We are using a static cut of the data, which has further been restricted to the period from January 2021 to December 2023 as this period had the greatest coverage for the subset.
Data processing of underlying firm data
Due to the noisy and complex nature of global supply chains and firm-level transaction data, we completed additional processing of the data to support our analysis. This involved converting historical transactions to use the most recent version of the Harmonised System for product classification (HS 2022), establishing firm ownership structures to reconcile subsidiaries at the country level, and addressing the set of firms that do not complete a transformation on the goods they import, or their imports are unrelated to their exports.
@@
As some transactions in our data use the previous 2017 Harmonised System of product classifications we convert these to the 2022 Harmonised System so that we have a consistent set. This was necessary as part of the period under study was in 2021 where the HS 2017 convention was still in place, and there were also cases where firms in later years had not updated their codes to the newer standard when completing customs declarations. We followed guidance produced by the Statistical Division of the UN for the conversion of products between the different harmonised systems which applies a matching process based on correlation tables.
@@
Within the data, transacting firms are reported at various levels within a company ownership structure; some may be wholly owned, while others might be a subsidiary of a multinational (or a subsidiary of a subsidiary … of a multinational).
The reconciliation of subsidiaries within country is important as the underlying data has instances whereby a subsidiary of a multi-national imports a product which is then used by a separate subsidiary of the same multi-national. In this case the imported goods are transferred between the subsidiaries domestically (under the same owner umbrella), but we do not capture these transfers as our data contains only cross-border trade.
@@
To rationalise the inconsistent ownership hierarchy, we use GSCIP’s ownership hierarchy dataset which sets out (where available) the global ultimate owner of subsidiaries. For each firm-entity we climb the ownership hierarchical structure until we get to the highest level of ownership up to the point of operation (as we do not want to classify distinct firms i.e. Heinz and Duracell into a larger non-distinct owner such as Berkshire and Hathaway.) To do this we stop climbing the ownership hierarchy to a higher-level firm when that firm can be classified as a financial corporation.
We also faced the issue of circular ownership - whereby a firm is owned by another one or more firms that are (at some tier) owned by the original firm - creating an infinite loop of ownership. To address this problem we break the circular ownership choosing one of the firms within the loop.
The ownership information we rely on is non-uniform and relationships can be opaque. This process improved the information we have on the firm set - but is however not a perfect process.
Once we have assigned subsidiaries to larger conglomerates we create a country-based version of each firm. So that, for example, a multinational firm that transacts to and/or from the UK would be treated as distinct to the same multinational that transacts to and/or from the US.
@@
Community detection. We present the number of communities, the Stability coefficient and the Variation of Information for the SC and AI networks for each partition in Figure S2. As the partition number increases, the algorithm runs for a longer ’time’, and larger (but typically fewer) communities are detected.
At each scale (partition number), the algorithm seeks to minimise the Stability coefficient and hence the communities are better defined for lower values. In addition, a low value of the Variation of Information tells us we find similar partitions for different random seeds. We observe that both the SC and AI network have very well defined communities at large partition numbers (low Stability and low Variation of Information), and hence we focus on communities in this region. In Figure S2 C-E, we visualise partitions 20, 25 and 30.
In Figure S3 we show two partitions of the NAFTA IO network, which exhibit similarities to the SC and AI networks.
@@
10 year growth regressions. In Figure S4 we replicate Figure 7 C and F for 10 year growth regressions (2011-2021). We find similar patterns with a slightly lower AUC in both cases.
Figure S4: 10 year growth regressions:

### v3

@@ Acknowledgements
This paper relies on data accessed through the UK HM Government’s Global Supply Chain Intelligence Programme (GSCIP). This contains firm-to-firm cross-border transactions, derived from customs and shipping data, which was compiled by a commercial provider. Further information on firms in the transaction dataset has been gathered by GSCIP from business registries, tax filings, company websites, and other sources.
The dataset includes ’primary’ customs data from a set of available country sources. All other countries are represented in the dataset as both customers and sellers of these primary countries. The exact set of primary countries included within the GSCIP programme is suppressed for reasons of sensitivity. We do, however, provide an overview of country-level coverage relative to total trade volumes in Figure 1.
While GSCIP is an ongoing programme, we use data from the period from January 2021 to December 2023, accessed in early 2025.
Processing of underlying firm data
Due to the noisy and complex nature of global supply chains and firm-level transaction data, we completed additional cleaning and processing of the raw data as detailed below.
@@
As some transactions in our data use the previous 2017 Harmonised System of product classifications we convert these to the 2022 Harmonised System so that we have a consistent set. This was necessary as our data includes 2021 when the HS 2017 convention was still in place, and there were also cases where firms in later years had not updated their codes to the newer standard when completing customs declarations. We followed concordance tables produced by the Statistical Division of the UN [56] for the conversion of products between the different harmonised systems.
@@
The reconciliation of subsidiaries within country is important as the underlying data has instances whereby a subsidiary of a multi-national imports a product which is then used by a separate subsidiary of the same multi-national. In this case the imported goods are transferred between the subsidiaries domestically (under the same owner umbrella), but we do not capture these transfers as our data contains only cross-border trade. For example, Cadberrys may import cocoa beans using one subsidiary and produce a chocolate bar using another subsidiary.
Transacting firms are reported at various levels within a company ownership structure; some may be wholly owned, while others might be a subsidiary of a multinational (or a subsidiary of a subsidiary … of a multinational). For each firm-entity we climb the ownership hierarchical structure, provided by GSCIP, until we get to the highest level of ownership up to the point of operation (as we do not want to classify distinct firms i.e. Heinz and Duracell into a larger non-distinct owner such as Berkshire Hathaway). To do this we stop climbing the ownership hierarchy when the next firm up is a financial entity or corporation.
Firm-level transaction network
Here we present descriptive statistics on the underlying inter-firm transaction data.
Degree statistics
Similar to other work to understand the structure of firm-level production networks [4], we analyse the in- and out-degree of our firm population.
The 1 billion transactions in our data correspond to 47 million firm-to-firm links. We consider this data to form a network in which nodes are firms and weighted edges correspond to the sum of the value of transactions between these firms within our time-period. The network is highly heterogeneous: the majority of firms (63%) maintain trading relationships with just one partner, while a minority engage with a large number of partners. We find that the maximum out-degree (number of export partners for a single firm) reaches 351,000, and the maximum in-degree (number of import partners) is 113,000. Notably, 1% of all firms have more than 50 trading partners, while only 0.01% have over 1,000 partners.
Within our full dataset of 47 million firms, we observe that 43% are exclusively exporters, 48% are exclusively importers, and only 9% both import and export. This latter set includes a core set of 1.8 million firms which are use to construct our production network (the remainder - exporter- and importer-only firms - are origins or destinations of transactions from this core set).
Firm types by geography
There are 8,895,000 firms (43% of all firms in the dataset) that export only (i.e., without any recorded imports). This pattern is especially strong in China (90.1% sender-only), Taiwan (81.1%), Switzerland (80.0%), Hong Kong (75.3%), Japan (72.8%), and South Korea (70.6%).
There are 9,771,000 nodes that only import (have no recorded exports). The representation of import-only nodes is most notable in Ecuador (88.8%), Peru (87%), the Philippines (81.5%), Canada (74.7%), and Ukraine (67.6%).
There are 1,803,000 firms (9% of all firms) that export and import within our dataset. These firms are especially prevalent in Vietnam (21%), Turkey (18%), the United Kingdom (17%), India (16%) and Germany (16%).
@@
Formulas
For our main analysis, we set stringent conditions to identify firms that both produce and purchase specific products:
We also produced alternative networks with variations to these criteria, including:
  1. 1.
    filtering the network to just include firms buying more of product $i$ than the average firm, without any conditions on selling product $j$.
  2. 2.
    filtering for just firms selling more of product $i$ than the average firm, without any conditions on buying product $i$.
  3. 3.
    Including all firms selling product $j$ and buying product $i$.
We replicate Figure 5 in the main text which displays edge, in-, and out-degree correlation with the ’AI network’ in Figure S1. We find that the conditions we apply in the main analysis result in stronger correlations with the AI network, regardless of thresholds on edge or firm count.
Community detection
We present the number of communities, the Stability coefficient and the Variation of Information for the SC and AI networks for each partition in Figure S3. As the partition number increases, the algorithm runs for a longer ’time’, and larger (but typically fewer) communities are detected.
At each scale (partition number), the algorithm seeks to minimise the Stability coefficient and hence the communities are better defined for lower values. In addition, a low value of the Variation of Information tells us we find similar partitions across a 1000 random seeds. We observe that both the SC and AI network have very well defined communities at large partition numbers (low Stability and low Variation of Information), and hence we focus on communities in this region. In Figure S3 C-E, we visualise partitions 20, 25 and 30.
In Figure S4 we show two partitions of the product network constructed based on NAFTA ’Rules of Origin’ [38] which exhibit similarities to the SC and AI networks.
Additional econometric regressions
In Figure S5 we replicate Figure 7 B-D and F-H for 10 year growth regressions (2011-2021), finding very similar results. In Figure S5, we vary the ’number of neighbours’ parameter in the density metric for high/low ECI countries, for both downstream and upstream models. Overall there is a small but perceptible peak in the AUF (no FE) around $k=40$ for most models, and so we use this value in the main paper. This is consistent with previous work [35] which found a peak of around $k=50$ in similar contexts.
Figures
Figure S1: Variations in the underlying formula that forms the product network: edge, in and out-degree correlations with the AI network. Figures show variation for setting Firm count and edge threshold at 1, 2, 3, 4 and 5. The correlation of the full edge set of AI network and SC network is highest with thresholds for both export and import.
@@
Figure S5: 10 year growth regressions: all countries, high ECI countries, low ECI countries.
Figure S6: Variation in density parameter k for high and low ECI country sets. We see the AUC peaks around k=40 for all cases (more darker squares).